Hi I have a substitution problem. I think I have the answer ,but I don't think I am doing it correctly.Can you tell me if I am correct and can you show me your way ,the easiest way to do this using the correct answer? Somewhere in the steps, I am messing up. This problem is a handout and not in a book. The problem is:
x=3y+7 and 2x+5y=3
The choice of answers are:
A) 4,1 B) 6,3 C) 1,-2 D)4,-1 I CHOSE D (4 ,-1)

Thankyou so much!

Answer by checkley71(8403)

X=3Y+7 NOW REPLACE X IN THE OTHER EQUATION BY (3Y+7) & SOLVE FOR Y.
2(3Y+7)+5Y=3
6Y+14+5Y=3
11Y=3-14
11Y=-11
Y=-11/11
Y=-1 ANSWER FOR Y.
X=3*-1+7
X=-3+7
X=4 ANSWER FOR X.
(4,-1) IS THE ANSWER D).